OpenStudy (anonymous):

work out problem: (3a^4)^3

OpenStudy (anonymous):

9a^12

OpenStudy (eyust707):

ok so we know that \[(3a^4)^3 = (3a^4)(3a^4)(3a^4) =(3^3)(a^{12}) \]

OpenStudy (eyust707):

\[=27a^{12}\]